# One Model For All: Partial Diffusion for Unified Try-On and Try-Off in Any Pose

### v1

@@ 3.2 Partial Diffusion
$\mathcal{L}_{\text{partial}}=\mathbb{E}_{z_{0},\epsilon,t,c}\left[\left\|M\cdot\left(\epsilon-\epsilon_{\theta}(\hat{z}_{t},c,t)\right)\right\|_{2}^{2}\right],$ (4)
where $\epsilon_{\theta}$ is the noise prediction network and $c$ is the conditional input. Components not selected for diffusion remain fixed throughout the process, allowing controllable and task-specific generation.
@@
LADI-VTON 0.856 0.137 12.072 5.763 0.821 0.810 15.925 7.732 7.52
StableGarment 0.827 0.115 12.463 6.311 0.805 0.799 18.539 10.508 7.59
StableVITON 0.849 0.123 9.566 3.062 0.819 0.800 12.489 4.124 7.56
OOTDiffusion 0.827 0.107 8.242 1.895 0.831 0.809 12.817 3.909 7.80
IDM-VTON 0.857 0.085 7.117 1.813 0.846 0.821 10.619 2.703 8.17
MV-VTON 0.879 0.114 7.980 2.760 0.841 0.827 11.002 2.906 8.07
CatVTON 0.843 0.099 7.074 1.890 0.833 0.817 12.178 3.987 7.98
@@
3.3 One Model for Try-on and Try-off
As shown in Fig. 2, our proposed OMFA framework supports both try-on stream and try-off stream, enabling remove garments from a source person and transfer onto a target person. To enable high-fidelity texture preservation, prior studies (Xu et al. 2024; Choi et al. 2024; Kim et al. 2024) introduce an additional ReferenceNet to extract garment features. However, this approach - where the main UNet extracts person features while ReferenceNet extracts garment features, with subsequent feature injection via attention mechanisms - exhibits inherent redundancy. In contrast, we concatenate multiple conditions of person $I_{p}$, garment $I_{g}$ and head images $I_{h}$ in the spatial dimensions and process them in a single UNet. The guiding SMPL-X pose $I_{s}$ image and the person image are both encoded into the latent space by the VAE encoder $\mathcal{E}$, and then concatenated since they are spatially aligned.
$z_{0}=\begin{cases}(\mathcal{E}(I_{p})\odot\mathcal{E}(I_{s}))\oplus\mathcal{E}(I_{g})\oplus I_{s}&\text{if Try-on}\\ (\mathcal{E}(I_{p})\odot\mathcal{E}(I_{s}))\oplus\mathcal{E}(I_{g})&\text{else Try-off}\end{cases},$ (5)
where $\odot$ and $\oplus$ indicate concatenation operation along the channel dimension and the spatial dimension respectively. Then the partial noise in Eq. 3 is added onto the latent $z_{0}$:
$z_{t}=\begin{cases}M_{1}\cdot z_{t}+(1-M_{1})\cdot z_{0}&\text{if Try-on}\\ M_{2}\cdot z_{t}+(1-M_{2})\cdot z_{0}&\text{else Try-off}\end{cases},$ (6)
where $M_{1}$ and $M_{2}$ denotes the mask for the person image $I_{p}$ and the garment image $I_{g}$. During inference, the diffusion step is repeated for $T$ times to predict the output latent $z^{\prime}_{0}$. Then, we split the output latent across the spatial dimension in width and transform the split latent with the VAE decoder $D$ to obtain the corresponding clothed person or garment.
$I_{output}=D(Split(z^{\prime}_{0}))$ (7)
3.4 SMPL-X-based Structural Conditioning
To introduce explicit and controllable human geometric information into the generation process, we propose a SMPL-X-based Structural Conditioning mechanism, in which the guiding image $I_{s}$ is used as the structural representation. The SMPL-X model is a low-dimensional parametric model of the human body that jointly use shape parameters $\beta\in\mathbb{R}^{10}$ and pose parameters $\theta\in\mathbb{R}^{24\times 3\times 3}$ to generate a 3D human mesh $M\in\mathbb{R}^{3\times N}$ with $N=6890$ vertices. To obtain the 3D information of the person image, we adopt an existing framework known as 4D-Humans (Goel et al. 2023) to regress the shape $\beta$, pose $\theta$ parameters from the person image, and use them to construct a 3D human mesh. We then apply the predicted camera parameters $\pi$ to render the mesh into an RGB image $I_{s}$. Using the camera projection function($\Pi$), the rendering process is formulated as:
$I_{s}=\Pi(\text{SMPL}(\beta,\theta),\pi)$ (8)
During the denoising process, we concatenate the latent $\mathcal{E}(I_{s})$ with the person latent $\mathcal{E}(I_{p})$ along the channel dimension.
A key advantage of using the SMPL-X model is its disentangled representation of pose and body shape, which enables pose-transfer try-on while preserving the body shape. By fixing the shape parameters and editing only the pose parameters, we are able to render maps of poses as the conditional input to guide try-on without additional template images.
4 Experiments
@@ 4.1 Experimental Setup
We train and evaluate our model on two publicly available fashion datasets: VITON-HD (Choi et al. 2021) and DeepFashion-MultiModal dataset (Liu et al. 2016; Jiang et al. 2022; Huang et al. 2024). VITON-HD contains 13,679 image pairs of frontal half-body models and corresponding upper-body garments, with 11,647 for training and 2,032 for testing. In the DeepFashion-MultiModal dataset, each sample includes not only images of person and garment but also a pair of target images in two poses. We select around 40,000 as training samples and 1,100 test samples. To prepare the inputs, we adopt SCHP (Li et al. 2020) to obtain semantic segmentation maps of different body regions.
@@
In our experiments, we initialize the models by inheriting the pretrained weights of Stable Diffusion XL, and finetune the parameters of the denoising UNet with the AdamW optimizer (Loshchilov and Hutter 2017), using $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model is trained at a high resolution of $768\times 1024$ on 4 NVIDIA A800 GPUs for 65,000 steps, with a batch size of 8 and a learning rate of $1e^{-6}$. To enable classifier-free guidance (Ho and Salimans 2022) and maintain generation diversity, we randomly drop each conditional reference feature with a probability of 0.05. During inference, we adopt the DDIM sampler (Song, Meng, and Ermon 2020) with 50 diffusion steps and set the guidance scale to 2.0.
@@
For the try-on task, we compare our method with the seven state-of-the-art methods : LADI-VTON (Morelli et al. 2023), StableGarment (Wang et al. 2024), StableVITON (Kim et al. 2023), OOTDiffusion (Xu et al. 2024), IDM-VTON (Choi et al. 2024), CatVTON (Chong et al. 2024) and MV-VTON (Wang et al. 2025). Under the realistic assumption that template garments are unavailable, we adapt the input pipelines of these methods by either: (1) employing segmented garment images, or (2) directly utilizing dressed person images. For the multi-pose try-on task, we compare our method with the baseline IDM-VTON (Choi et al. 2024). For the try-off task, we evaluate our method against two recent approaches: TryoffDiff (Velioglu et al. 2024) and TryoffAnyone (Xarchakos and Koukopoulos 2024). We employ pre-trained checkpoints provided in official repositories for comparison.
@@
For paired settings, we adopt four widely used metrics to evaluate the similarity between synthesized images and ground-truth images: SSIM (Wang et al. 2004), LPIPS (Zhang et al. 2018), FID (Heusel et al. 2017), and KID (Bińkowski et al. 2018). For unpaired settings, in addition to computing FID and KID, we further calculate CLIP-I (Radford et al. 2021) and DINO (Oquab et al. 2023) similarity between the segmented garment region and the corresponding reference garment to assess garment-level semantic consistency. To ensure a fair comparison with mask-based methods, our method utilizes the agnostic map to maintain the unedited regions unchanged, similar to CatVTON (Chong et al. 2024). Moreover, given the person and garment images, we use GPT-4o-mini to provide a comprehensive score for the try-on result. The score ranges from 0 to 10. For the garment generation task, we additionally report DISTS (Ding et al. 2020), a perceptual similarity metric designed to capture both structural and textural fidelity between the generated garment image and the ground truth.
@@ 4.2 Virtual Try-on
Tab. 1 presents the virtual try-on results on the VITON-HD dataset. In the paired setting, our method achieves comparable overall metrics. While some baselines report slightly higher SSIM scores, this can be attributed to the fact that the input warped cloth is already well-aligned with the target, making it easier for models to preserve realistic garment appearance. Benefiting from the ability to reconstruct garments, our method performs notably better in the more challenging unpaired try-on setting, particularly in terms of the CLIP-I and DINO similarity. In Fig. 3, we present a qualitative comparison between our method and multiple advanced approaches on the VITON-HD dataset, highlighting its distinct advantage in the person-to-person try-on scenario. Specifically, existing methods lack effective generalization under garment occlusion, deformation, or pose misalignment, often leading to texture loss, artifacts, or contour distortion. In contrast, our mask-free and unified framework for try-off and try-on is capable of producing high-fidelity results.
@@ 4.3 Virtual Try-off
Tab. 3 shows the quantitative comparison of virtual try-off. Our method outperforms advanced approaches across all five metrics, showing significant advantages in detail preservation, structural and textural consistency, and semantic alignment. Fig. 4 presents a comparison of garment reconstruction results on the VITON-HD dataset between our method and other try-off approaches. Specifically, TryOffDiff tends to capture only the coarse structure and color of the garment, failing to restore fine-grained details of the patterns. While TryOffAnyone is capable of handling complex and irregular patterns, it still suffers from blurred or missing text elements. By comparison, our method shows clear and consistent advantages in detail preservation, particularly in the clarity of textual contours and pattern boundaries.
@@
Figure 6: Visual comparisons of our proposed modules. We adopt the parallel UNet architecture as the baseline, and introduce partial diffusion and a single UNet for processing joint inputs. “One Model For All” indicates unifying try-off and try-on tasks in a single UNet.
Methods Paired Unpaired
@@
Parallel UNets 0.839 0.160 19.139 9.841 0.805
Single UNet 0.853 0.098 12.663 3.875 0.860
One Model For All 0.862 0.098 10.527 1.920 0.876
Table 4: Ablation study of the proposed modules on the VITON-HD dataset. We use the parallel UNet architecture as the baseline. Incorporating partial diffusion and replacing the parallel UNets with a single UNet for joint input processing leads to consistent improvements across all evaluation metrics and reduces computational cost. Our OMFA configuration unifies try-on and try-off tasks within a single UNet, further enhancing garment transfer performance.
4.4 Ablation Studies
We perform ablation studies to validate the effectiveness of the proposed partial diffusion mechanism and the unified generation strategy for try-on and try-off in the person-to-person try-on scenarios. In the baseline setting, we train parallel UNets following IDM-VTON (Choi et al. 2024), which utilizes a ReferenceNet to encode the garment images and inject the garment features into the denoising UNet. As shown in Fig. 6, this approach often leads to texture loss and inaccurate garment structures when using the warped garment as input. We then replace ReferenceNet with a single UNet and use the proposed partial diffusion mechanism to handle the spatially joint input, which effectively improves the performance and decreases the computation cost. Finally, we unified the try-on and the try-off tasks using a single UNet, which is noted as “One Model For All”, resulting in clearer texture restoration and more complete garment contours. The quantitative results in Tab. 4 further validate the effectiveness of our proposed method.
@@ 5 Conclusion
We presented OMFA, a unified diffusion-based framework for virtual try-on and try-off that overcomes key limitations of prior methods, including dependence on garment templates, segmentation masks, and fixed poses. OMFA introduces a novel partial diffusion mechanism for efficient, bidirectional garment-person transformation with fine-grained subtask control. It operates in a fully mask-free manner and requires only a single portrait and target pose, making it practical for real-world use. With SMPL-X–based pose conditioning, OMFA enables flexible, multi-view try-on from a single image. Extensive experiments confirm its effectiveness and generalizability across both tasks.
Appendix A Implementation Details
A.1 Training and Inference Details
We first train the model on the VITON-HD dataset with a resolution of $384\times 512$ for 20K iterations. Then, keeping the same learning rate and batch size, we finetune the model on both the VITON-HD and DeepFashion-MultiModal datasets using a higher resolution of $768\times 1024$. For data augmentation, we enhance the background color of the generated garment. Specifically, we use a tensor of the same size as the input garment, with all values set to 255, and concatenate it with the garment latent $\mathcal{E}(I_{g})$ along the channel dimension. To align the latent with the UNet input along the channel dimension, we apply separate convolutional layers to each component of the joint input, projecting their channels to 320. Each convolutional layer is initialized with the first several channels of the corresponding layer in the pretrained UNet. During inference, if the try-on task is required, the inputs needed are the the person image $I_{p}$, the garment $I_{g}$, and the person’s portrait $I_{h}$, whereas for the try-off task, only $I_{p}$ is needed, with the other inputs set to 0. Our implementation is based on the PyTorch deep learning framework (version 2.1.2), with the diffusion model adapted from HuggingFace’s Diffusers library.
A.2 Evaluation Metrics
In our experiments, we adopt a variety of evaluation metrics commonly used in generative tasks. Among them, SSIM (Wang et al. 2004), LPIPS (Zhang et al. 2018), KID (Bińkowski et al. 2018), and FID (Heusel et al. 2017) are widely used general metrics in related work. This section focuses on the detailed computation of several additional quantitative metrics used in our method, including DINO similarity (Oquab et al. 2023), CLIP-I (Radford et al. 2021), and LLM-based Image Scoring.

### v3

@@ 3.3 Bidirectional Tweedie Diffusion
$\mathcal{L}_{\mathrm{diff}}=\mathbb{E}_{m,t,\epsilon}\left[\left|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t}^{m},t,\mathcal{C}^{m})\right|_{2}^{2}\right].$ (8)
By minimizing Eq. (8), the network intrinsically converges to the conditional expectation required by Eq. (7). Thus, optimizing the standard noise prediction loss is not an arbitrary choice, but the mathematically exact implementation required to evaluate Tweedie’s sequence-recovery formula. At inference, the target latent is iteratively denoised under the fixed condition set $\mathcal{C}^{m}$ using the network $\epsilon_{\theta}$. The final recovered latent is then decoded to image space:
$\hat{I}_{d}=\mathcal{D}(\hat{\mathbf{z}}_{0}^{d}),\qquad\hat{I}_{g}=\mathcal{D}(\hat{\mathbf{z}}_{0}^{g}).$ (9)
3.4 SMPL-X-based Structural Conditioning
To introduce explicit and controllable human geometric information into the generation process, we propose a SMPL-X-based structural conditioning mechanism, where the rendered structural image $I_{s}$ serves as the geometric condition. SMPL-X is a low-dimensional parametric human body model that jointly uses shape parameters $\beta\in\mathbb{R}^{10}$ and pose parameters $\theta\in\mathbb{R}^{24\times 3\times 3}$ to produce a 3D human mesh with $N=6890$ vertices. Given a person image, we use 4D-Humans [10] to regress the corresponding shape parameters $\beta$, pose parameters $\theta$, and camera parameters $\pi$. The estimated SMPL-X mesh is then rendered into an RGB structural image $I_{s}$ through the camera projection function $\Pi$, formulated as
$I_{s}=\Pi(\text{SMPL-X}(\beta,\theta),\pi).$ (10)
The rendered structural image is further encoded into a structural latent:
$\mathbf{z}_{s}=\mathcal{E}(I_{s}).$ (11)
During denoising, we concatenate the structural latent $\mathcal{E}(I_{s})$ with the person latent $\mathcal{E}(I_{p})$ along the channel dimension to obtain a structure-aware person latent $\tilde{\mathbf{z}}_{p}$.
A key advantage of SMPL-X lies in its disentangled representation of body shape and pose. By fixing the shape parameters while editing only the pose parameters, we can render structural conditions under arbitrary target poses while preserving the body shape of the source person. This enables pose-controllable virtual try-on without requiring additional template images.
4 Experiments
@@
LADI-VTON [27] 0.856 0.137 12.072 5.763 0.821 0.810 15.925 7.732 7.52
StableGarment [40] 0.827 0.115 12.463 6.311 0.805 0.799 18.539 10.508 7.59
StableVITON [21] 0.849 0.123 9.566 3.062 0.819 0.800 12.489 4.124 7.56
OOTDiffusion [46] 0.827 0.107 8.242 1.895 0.831 0.809 12.817 3.909 7.80
IDM-VTON [5] 0.857 0.085 7.117 1.813 0.846 0.821 10.619 2.703 8.17
MV-VTON [39] 0.879 0.114 7.980 2.760 0.841 0.827 11.002 2.906 8.07
CatVTON [6] 0.843 0.099 7.074 1.890 0.833 0.817 12.178 3.987 7.98
CatVTON [6]+TryOffDiff [36] 0.836 0.118 12.611 3.663 0.825 0.821 12.611 3.663 -
IDM-VTON [5]+TryOffDiff [36] 0.859 0.090 7.849 1.312 0.847 0.830 11.299 2.670 -
MV-VTON [39]+TryOffDiff [36] 0.879 0.114 8.082 2.857 0.840 0.827 11.174 3.288 -
Any2AnyTryon [12] 0.861 0.151 7.168 1.732 0.829 0.818 13.398 5.401 -
@@ 4.1 Experimental Setup
We train and evaluate our model on two publicly available fashion datasets: VITON-HD [4] and DeepFashion-MultiModal dataset [25, 20, 17]. VITON-HD contains 13,679 image pairs of frontal half-body models and corresponding upper-body garments, with 11,647 for training and 2,032 for testing. In the DeepFashion-MultiModal dataset, each sample includes not only images of person and garment but also a pair of target images in two poses. We select around 40,000 as training samples and 1,100 test samples. To prepare the inputs, we use a segmentation method [23] to obtain different image regions.
@@
In our experiments, we initialize the models by inheriting the pretrained weights of Stable Diffusion XL, and finetune the parameters of the denoising UNet with the AdamW optimizer [26], using $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model is trained at a high resolution of $768\times 1024$ on 4 NVIDIA A800 GPUs for 65,000 steps, with a batch size of 8 and a learning rate of $1e^{-6}$. To enable classifier-free guidance [16] and maintain generation diversity, we randomly drop each conditional reference feature with a probability of 0.05. During inference, we adopt the DDIM sampler [35] with 50 diffusion steps and set the guidance scale to 2.0.
@@
For the try-on task, we compare our method with several state-of-the-art methods, including LADI-VTON [27], StableGarment [40], StableVITON [21], OOTDiffusion [46], IDM-VTON [5], CatVTON [6] and MV-VTON [39]. Under the more realistic setting where template garments are unavailable, we adapt the input pipelines of these methods to use segmented garment images. For the multi-pose try-on task, we compare our method with the baseline IDM-VTON [5]. For the try-off task, we evaluate our method against two recent approaches: TryOffDiff [36] and TryoffAnyone [42]. Additionally, to evaluate cross-task compatibility, we consider five two-stage pipelines that perform garment synthesis followed by try-on. Our method and Any2AnyTryon [12] both follow this two-stage design. We also include three additional combinations that feed TryOnDiff-generated garments into CatVTON [6], IDM-VTON [5], and MV-VTON [39]. We use the pre-trained checkpoints provided in the official repositories of all compared methods.
@@
For paired settings, we evaluate similarity between synthesized and ground-truth images using SSIM [41], LPIPS [51], FID [15], and KID [1]. For unpaired settings, in addition to computing FID and KID, we further calculate CLIP-I [32] and DINO [29] similarity between the segmented garment region and the corresponding reference garment to assess garment-level semantic consistency. To ensure a fair comparison with mask-based methods, we use the agnostic map to preserve the unedited regions, following CatVTON [6]. Moreover, given the person and garment images, we use GPT-4o-mini to provide a comprehensive score for the try-on result. The score ranges from 0 to 10. For the garment generation task, we additionally report DISTS [8], a perceptual similarity metric designed to capture both structural and textural fidelity between the generated garment image and the ground truth.
Figure 5: Qualitative comparison of TryOffDiff-combined try-on pipelines and our unified framework. Methods combined with TryOffDiff tend to blur patterns, whereas our method better preserves garment details.
@@ 4.2 Virtual Try-on
Tab. 1 reports the virtual try-on results on the VITON-HD dataset. In the paired setting, our method achieves comparable overall metrics. While some baselines report slightly higher SSIM scores, this is most likely due to the input warped cloth being well aligned with the target, making it easier to maintain garment appearance. When garments recovered by existing try-off methods are used as inputs to the try-on framework (see Fig. 5 and Fig. 6), fine-grained texture details remain inadequately recovered. Although such inputs usually provide a complete garment silhouette, the subsequent try-on stage still exhibits degraded fabric realism or blurred patterns/LOGOs. Benefiting from the ability to reconstruct garments, our method performs notably better in the more challenging unpaired try-on setting, particularly on CLIP-I and DINO similarity. In Fig. 3, we present a qualitative comparison between our method and multiple advanced approaches on the VITON-HD dataset, highlighting its distinct advantage in the person-to-person try-on scenario.
Multi-pose Virtual Try-on.
@@ 4.3 Virtual Try-off
Tab. 3 shows quantitative results for virtual try-off, where our method outperforms advanced approaches across all five metrics, showing significant advantages in detail preservation, structural and textural consistency, and semantic alignment. Fig. 6 presents a comparison of garment reconstruction results on the VITON-HD dataset between our method and other try-off approaches. Specifically, TryOffDiff mainly recovers coarse shape and color and often misses fine patterns, while TryOffAnyone better handles complex patterns but still blurs or omits text. By comparison, our method shows clear and consistent advantages in detail preservation, particularly in the clarity of textual contours and pattern boundaries.
@@
4.4 Ablation Studies
Effectiveness of the Bidirectional Tweedie Diffusion.
In the baseline setting, we follow IDM-VTON [5] and train parallel UNets with a ReferenceNet that encodes the garment images and injects their features into the denoising UNet. We then replace ReferenceNet with a single UNet and adopt the Bidirectional Tweedie Diffusion to process the spatially joint input. As shown in Fig. 7(c), even with a warped garment input, this mechanism delivers more faithful details compared to (a). The quantitative results in the first group (lines 1 and 3) and the second group (lines 2 and 4) of Tab. 4 further demonstrate improved visual fidelity.
Table 4: Ablation study of the proposed modules on the VITON-HD dataset. We use the Dual UNet architecture as the baseline, where “Single UNet” denotes our LLM-inspired bidirectional diffusion design. Our OMFA achieves consistently superior scores on all evaluation metrics.
Arch. Garment Input Paired Unpaired
@@ 4.4 Ablation Studies
Dual UNet Segmented 0.839 0.160 19.139 9.841 0.805
Dual UNet Try-off 0.843 0.148 13.587 4.707 0.856
Single UNet Segmented 0.853 0.098 12.663 3.875 0.860
Single UNet Try-off 0.862 0.098 10.527 1.920 0.876
Effectiveness of the unified generation strategy.
We further validate the unified generation strategy for try-on and try-off in person-to-person scenarios. When the exemplar garment is unavailable, try-on results with segmented-garment input may exhibit incomplete contours and occluded details (see Fig. 7(a) and (c)). In contrast, our unified pipeline first performs try-off and then conducts garment transfer within a single model, producing the most complete and visually consistent results . The quantitative results in the two ablation groups (lines 1 and 2, and lines 3 and 4 in Tab. 4) further support this conclusion.
Figure 7: Visual comparisons of our proposed modules. We also compare variants using a segmented input versus a try-off input.
Impact of SMPL-X-based conditioning.
We additonally train two small-resolution variants on VITON-HD, respectively using DensePose and SMPL-X as the body-structure inputs. As shown in Tab. 5, SMPL-X performs similarly to DensePose. However, as illustrated in Fig. 8, DensePose sometimes causes shape distortion (e.g., arms looking thinner; left) and boundary artifacts (right). By contrast, SMPL-X offers explicit geometric modeling, allowing direct control of body pose and shape through its parameters.
Figure 8: Qualitative ablation of structural conditioning on VITON-HD, illustrating the benefit of SMPL-X over DensePose.
Table 5: Quantitative ablation of structural conditioning on VITON-HD: DensePose vs. SMPL-X.
[…]
Densepose 0.839 0.132 0.863 11.645 3.610
SMPL-X 0.839 0.152 0.872 11.781 3.602
@@ 5 Conclusion
We presented OMFA, a unified diffusion-based framework for virtual try-on and try-off that overcomes key limitations of prior methods, including dependence on garment templates, segmentation masks, and fixed poses. OMFA introduces a novel bidirectional Tweedie diffusion mechanism for efficient, interactive garment-person transformation with fine-grained subtask control. It operates in a fully mask-free manner and requires only a single portrait and a target pose, making it practical for real-world use. With SMPL-X–based pose conditioning, OMFA enables flexible, multi-view try-on from a single image. Extensive experiments confirm its effectiveness and generalizability across both tasks.
6 Acknowledgments
This work was supported in part by National High-Level Young Talent Program (Grant 2025HY00260104), in part by the Fundamental Research Funds for Higher Education Institutions allocated to Sun Yat-sen University (Grants 25hytd007 and 2025RGZN009), in part by the Guangdong Provincial High-Level Young Talent Program (Grant 2025HYSPT0707), in part by the Tuoyuna Grant (HT-99982025-0564), in part by the Faculty Start-up Research Fund (Grant 67000-12255002), and in part by the Huawei Strategic Research Institute Talent Fund.
Supplementary Material
7 Implementation Details
7.1 Training and Inference Details
We first train the model on the VITON-HD dataset with a resolution of $384\times 512$ for 20K iterations. Then, keeping the same learning rate and batch size, we finetune the model on both the VITON-HD and DeepFashion-MultiModal datasets using a higher resolution of $768\times 1024$. For data augmentation, we enhance the background color of the generated garment. Specifically, we use a tensor of the same size as the input garment, with all values set to 255, and concatenate it with the garment latent $\mathcal{E}(I_{g})$ along the channel dimension. To align the latent with the UNet input along the channel dimension, we apply separate convolutional layers to each component of the joint input, projecting their channels to 320. Each convolutional layer is initialized with the first several channels of the corresponding layer in the pretrained UNet. During inference, if the try-on task is required, the inputs needed are the person image $I_{p}$, the garment $I_{g}$, and the person’s portrait $I_{h}$, whereas for the try-off task, only $I_{p}$ is needed, with the other inputs set to 0. Our implementation is based on the PyTorch deep learning framework (version 2.1.2), with the diffusion model adapted from HuggingFace’s Diffusers library.
7.2 Evaluation Metrics
In our experiments, we adopt a variety of evaluation metrics commonly used in generative tasks. Among them, SSIM [41], LPIPS [51], KID [1], and FID [15] are widely used general metrics in related work. This section focuses on the detailed computation of several additional quantitative metrics used in our method, including DINO similarity [29], CLIP-I [32], and LLM-based Image Scoring.